Run Planning Meeting Notes, April 17 - April 23, 2025
From GlueXWiki
↑ Back to Run Coordination Meetings:Spring2025#Runplan
Contents
Tuesday April 22
Shift Crew Summaries
- Day 0800-1600 (Nice template developed by Kevin Scheuer)
- Evening 1600-2400
- Owl April 23, 0000-0800
Reports
- ECAL missing \pi^0 problem: solved?
- ECAL HV adjusted HallD Run #131151: Ecal scaler rates
- Updated a few rootspy references with run 131153
- new RootSpy plot added: ECAL-RF timing
Issues
Plans
- Maintenance day:
- SAF113 training (8+1 person)
- Radcon survey for restrictive access
- parasitic experiments
- EIC + MOLLER (how many people?)
- open the Hall gate
- Dissasemble parts, Crane use
- need to wrap up by 3-4pm
Monday April 21
Shift Crew Summaries
Reports
Issues
- DAQ RocTrig2 failed again.
- Empty scalar readout channels in TAGH and PS since weekend status: TAGH scalers status: PS scalers
- DIRC scaler
- TOF ADC North 41 hot channel
- FCAL hot channel
Plans
- Diamond Alignment when beam is good
- FCAL trigger with new voltage (start from midnight, run 131133)
- Prepare for maintanence day tomorrow (Wed):
- Restricted access for parasitic experiments
- Crane use
- SAF113 training
Weekend April 19-20
Shift Crew Summaries
- April 18: Day (0800-1600) Evening (1600-2400)
- April 19: Owl (0000-0800) Day (0800-1600) Evening (1600-2400)
- April 20: Owl (0000-0800) Day (0800-1600) Evening (1600-2400)
- April 21: Owl (0000-0800)
Reports
Sunday
Saturday
Friday
- FCAL only trigger setup for the overnight production running
- Comparision between ECAL+FCAL/BCAL trigger with and without the 3 ECAL layers turn off
- Comparison between ECAL/BCAL and ECAL+FCAL/BCAL trigger
- Thresholds determined for Start Counter discriminators
Issues
Sunday
Saturday
Friday
- Main trigger rate in EPICS is flat since about 11:46am
- Hall D beam energy not measured (IBCAD00CRCUR6) follow-up
- CAEN HV chassis causing problems
Plans
- ECAL bad channel testing
- Diamond Alignment: Hall A Reconfig (Tuesday-Thursday), start from Friday(4/25) 20 μA --> Monday(4/28) 50 μA
Friday April 18
Shift Crew Summaries
- 04/17 Owl 0000-0800
- 04/17 Day 0800-1600
- Around 13:30, recovered from 5th pass separator trips
- Harp Scan: Scan resultExtrapolation fit result
- 04/17 Evening 1600-2400
- 04/18 Owl 0000-0800
Reports
- Trigger for pi0 calibration
- Updated config's name: FCAL_BCAL_PS_2025_mod_p9.conf
- running it overnight (start from 131056 with 120 nA, now 131065), will check the pi0 yield
- Sean: updated the timing alignment calibrations as of run 131057
- Adjusting trigger timing offsets
- ECAL/FCAL energy calibration status
- TRD status: running with (ECAL+FCAL)+BCAL trigger
- 2025 MPGD Test beam setup
Issues
- Hall D Convergence
- TAGH: Cumulative stress on low-energy TAGH counters from persistent high occupancy at large beam current
- Example Run 131046 with 300nA
- Instruction on switching off low-energy TAGH counters
- TAGH mapping issues with two modules: 21 and 50
- CDC: New dark straw
- TOF readout window missing triggers
- FCAL base locked up after 131062 stopped
Plans
Thursday April 17
Last 24hr
Shift Crew Summaries
- Owl (April 16 0000-0800)
- Day (April 16 0800-1600)
- Evening (April 16 1600-2400)
- Harp scan: 5C11 5C11B radiatorHarp Scan Extrapolation
- Owl (Aprl-17 0000-0800)
- Harp scan:scan resultHarp Scan Extrapolation (compare with Last week, 04/09)
Reports
During RF Recovery Day Hall D Access
- TAGH CAEN HV Crate T1-8-TOP replaced.
- Related entry: TAGH Frozen Voltage Alarms: Channels 1-120, reported Tue, 04/15/2025
- Remote power supply reset verified; CPU reset worked on second attempt.
- Goniometer controller reset done
- ECAL module check
- Solenoid PXI Watchdog ramp down behavior disabled
- Related entry: Solenoid ramped down to zero, reported on Tue, 04/15/2025
- Added filter to barometric pressure
- Related entry: Noisy pressure signal, reported Tue, 04/15/2025
- Naomi's follow up
- TRD New Kr bottle
- FCAL base changes
Studies
Software
Issues
Plans
- Short term run plan
- highest priority: ECAL+FCAL+BCAL trigger preparing for pi0 production
- Diamond alignment
- Production: pi0 running
- anything else?